OpenStudy (anonymous):

3(x + 1)= -2(x- 1)- 4

OpenStudy (anonymous):

are you solving for x?

OpenStudy (shamil98):

3x + 3 = -2x + 2 - 4 5x= -5 x = -1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432 -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 1 points) (02.01) The triangle below has a total perimeter of 450 cm2. Write an equation and solve it to find the length of the longer side of the rectangle. 128x + 150 = 450; x = 2.34 cm x + 278 = 450; x = 172 cm 2x + 278 = 450; x = 86 cm 150x + 128 = 450; x = 2.15 cm -------------------------------------------------------------------------------- Question 5 (Multiple Choice Worth 1 points) (02.01) Solve for x: = 3x x = -7 x = -13 x = 8 x = 2 -------------------------------------------------------------------------------- Question 6 (Multiple Choice Worth 2 points) (02.01) Connie has saved up $15 to purchase a new CD from the local store. The sales tax in her county is 5% of the sticker price. Write an equation and solve it to determine the value of the highest priced CD Connie can purchase with her $15, including the sales tax. Round your answer to the nearest penny. x + 0.05x = 15; x = $14.29 x - 0.05x = 15; x = $15.79 0.5x = 15; x = $30 5x = 15; x = $3.00 -------------------------------------------------------------------------------- Question 7 (Fill-In-The-Blank Worth 2 points) (02.01) Solve for x: x + = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar. Answer for Blank 1: -------------------------------------------------------------------------------- Question 8 (Multiple Choice Worth 1 points) (02.01) Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution